amy needs to male 10 L of 42% sulphuric acid solution. in the supply room, she finds bottles of 30% sulphuric acid solution and 60% sulphuric acid solution. what volume of each solution should she mix in order to make the 42% solution?
Answer by richwmiller(17219)


By substitution
We know the total amount is 10 liters
We know the percentage of the total amount is 42%
We know the total amount has 4.2 liters of 100%
We want to know the amount of solution at 30%
We want to know the amount of solution at 60%
x+y=10,
0.3*x+0.6*y=0.42*10
0.3*x+0.6*y=4.2
x=10-y
0.3*(10-y)+0.6*y=4.2
3-0.3y+0.6*y=4.2
0.3*y=1.2
We now know amount of solution at 60%
y=4.0 liters at 60%
x=10-y
We now know amount of solution at 30%
x=6.0 liters at 30%
check
0.3*6.0+0.6*4.0=0.42*10
1.8+2.4=4.2
4.2=4.2
ok